# CAPE: A CLIP-Aware Pointing Ensemble of Complementary Heatmap Cues for Embodied Reference Understanding

### v1

@@ C.4 CLIP Fusion
In this method, we leverage both the CLIP similarity score and the model’s confidence score to make the final prediction. Specifically, we compute a hybrid score for each bounding box by summing its confidence score and its CLIP similarity score. We use the Top-2 predictions from both models for this process. However, the CLIP similarity scores and model confidence scores operate on different scales: while confidence scores are normalized in the range , CLIP similarity scores typically range around . To balance their influence and prevent the CLIP score from dominating, we scale the CLIP similarity scores by a factor of . The scaled CLIP score is then added to the corresponding confidence score to compute the final hybrid score for each bounding box. Finally, we select the bounding box with the highest hybrid score as the final prediction. The algorithm for this method is presented in Algorithm C.4.
C.5 CLIP-Aware Pointing Ensemble (CAPE)
@@
In this method, we combine the CLIP-Only (Top-2 + Threshold) and CLIP Fusion approaches. Through empirical analysis on validation set, we observe that for small objects, the CLIP Fusion method yields more accurate results than other ensemble strategies. This is primarily because CLIP tends to perform less reliably on smaller objects. Therefore, we apply CLIP Fusion to not rely on only CLIP-score but also consider the confidence score. To determine whether an object is small, we follow the definition provided by [11]: if the object occupies less than of the total image area, we classify it as small and use the CLIP Fusion ensemble. For all other cases, we apply the CLIP-Only (Top-2 + Threshold) method. The algorithm is presented in Algorithm C.5.
@@
Appendix D Finetuning or Frozen Text Encoder?
Based on multiple training experiments, we observed that finetuning the text encoder sometimes leads to instability, represented by the pink line in Fig. E.5. In contrast, freezing the text encoder results in consistently stable training, as shown by the green and orange lines in Fig. E.5. Therefore, achieving stable training with a finetuned text encoder requires several attempts. However, despite this challenge, finetuning the text encoder ultimately yields better performance.
Appendix E Hyperparameters and Other Training Parameters
In Tab. C.8, we present empirically determined hyperparameters and other key training details to support reproducibility.
Appendix F More Results
In Tab. F.9, we report the mAP at three different thresholds for all ablation setups, categorized by object size. Similarly, in Tab. F.10, we present the CLIP scores and metrics for the same setups, also with respect to object size.
In Fig. F.6, we present examples of failure cases. In the first row, although our model correctly detects the target object, CAPE selects the prediction from , which is incorrect in this case. In the second row, the bottles are extremely difficult to distinguish, even with explicit pointing cues. Nevertheless, the predicted bounding boxes cover other bottles adjacent to the GT one, which is acceptable, as they overlap with the pointing line and are semantically consistent with the text. In the third row, both of our models, as well as the Touch-Line models, produce incorrect predictions. However, guided by the pointing information, the predictions from and lie on the pointing line and are very close to the GT object. In contrast, the Touch-Line predictions are entirely off the pointing line and significantly distant from the GT object.
@@ Appendix F More Results
| Touch-Line-EWL [35] | - | - | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 | |
| Touch-Line-VTL [35] | - | - | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 | |
| 1 | Baseline | - | - | 0.7409 | 0.7003 | 0.7341 | 0.7861 | 0.2662 | 0.2182 | 0.2506 | 0.3266 |
| 2 | Setup 1 + object center prediction | - | - | 0.7339 | 0.6947 | 0.7277 | 0.7787 | 0.2707 | 0.2201 | 0.2692 | 0.3222 |
| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 0.7353 | 0.6934 | 0.7297 | 0.7812 | 0.2984 | 0.2763 | 0.2616 | 0.3533 |
| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 0.7375 | 0.6936 | 0.7333 | 0.7852 | 0.2694 | 0.2649 | 0.2233 | 0.3172 |
| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 0.7344 | 0.6939 | 0.7306 | 0.7780 | 0.2770 | 0.2568 | 0.2319 | 0.3382 |
| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 0.7373 | 0.6953 | 0.7343 | 0.7819 | 0.249 | 0.222 | 0.2202 | 0.3024 |
| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 0.7345 | 0.6965 | 0.7277 | 0.7780 | 0.2744 | 0.2540 | 0.2314 | 0.3335 |
| 8 | Setup 6 | Channel-wise input | - | 0.7348 | 0.6933 | 0.7273 | 0.7828 | 0.2959 | 0.2437 | 0.2642 | 0.3767 |
| 9 | Setup 6 | Channel-wise feature | - | 0.7372 | 0.6953 | 0.7311 | 0.7839 | 0.2598 | 0.2289 | 0.2119 | 0.3339 |
| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 0.7735 | 0.7441 | 0.7843 | 0.7923 | 0.247 | 0.2176 | 0.2042 | 0.3141 |
| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 0.7932 | 0.7688 | 0.8010 | 0.8098 | 0.2391 | 0.2242 | 0.2011 | 0.2878 |
| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 0.7379 | 0.7967 | 0.7343 | 0.7819 | 0.2422 | 0.2158 | 0.2009 | 0.3060 |
| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 0.7742 | 0.7927 | 0.7455 | 0.7842 | 0.2434 | 0.2137 | 0.2028 | 0.3090 |
| 14 | Setup 5 + Setup 6 (Final Model) | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 0.7382 | 0.6954 | 0.7323 | 0.7855 | 0.2634 | 0.2626 | 0.2233 | 0.3006 |
| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7387 | 0.6962 | 0.7325 | 0.7860 | 0.2695 | 0.2685 | 0.2344 | 0.3028 |
| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7391 | 0.6970 | 0.7332 | 0.7863 | 0.2765 | 0.2750 | 0.2499 | 0.3026 |
| Ours - Final | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 | |

### v4

@@ B.4 CLIP Fusion
In this method, we leverage both the CLIP similarity score and the model’s confidence score to make the final prediction. Specifically, we compute a hybrid score for each bounding box by summing its confidence score and its CLIP similarity score. We use the Top-2 predictions from both models for this process. However, the CLIP similarity scores and model confidence scores operate on different scales: while confidence scores are normalized in the range , CLIP similarity scores typically range around . To balance their influence and prevent the CLIP score from dominating, we scale the CLIP similarity scores by a factor of . The scaled CLIP score is then added to the corresponding confidence score to compute the final hybrid score for each bounding box. Finally, we select the bounding box with the highest hybrid score as the final prediction. The algorithm for this method is presented in Algorithm B.4.
B.5 CLIP-Aware Pointing Ensemble (CAPE)
@@
In this method, we combine the CLIP-Only (Top-2 + Threshold) and CLIP Fusion approaches. Through empirical analysis on validation set, we observe that for small objects, the CLIP Fusion method yields more accurate results than other ensemble strategies. This is primarily because CLIP tends to perform less reliably on smaller objects. Therefore, we apply CLIP Fusion to not rely on only CLIP-score but also consider the confidence score. To determine whether an object is small, we follow the definition provided by [13]: if the object occupies less than of the total image area, we classify it as small and use the CLIP Fusion ensemble. For all other cases, we apply the CLIP-Only (Top-2 + Threshold) method. The algorithm is presented in Algorithm B.5.
Appendix C Discussion
C.1 Why is pointing heatmap useful?
In ERU, the model must associate a textual instruction with the spatial region indicated by the person in the image. Without explicit spatial guidance, the model relies on implicit cues, such as hand position or object features, which become ambiguous when multiple candidate objects fit the text instruction. A pointing line provides a clear directional prior, narrowing the possible targets to regions intersecting or near the line. Formally, if is the set of candidate regions without explicit pointing heatmap (), then pointing guidance:
where denotes the image, refers to the textual description, and is the pointing line. This reduces ambiguity and helps the model converge to the correct region.
A second benefit is spatial alignment of multimodal embeddings. Without spatial cues, the latent representation of the intended object may blend with nearby distractors. The pointing line acts as a spatial attention prior, biasing the model toward regions consistent with the pointing direction.
Finally, using a pointing heatmap also strengthens the gradient signal during training. The heatmap provides an additional, well-structured supervisory cue, reducing overfitting to spurious text or background correlations. Effectively, serves as a spatial inductive bias that stabilizes and accelerates learning.
In summary, the pointing line provides an explicit spatial prior, constrains the candidate space, improves multimodal alignment, and yields more efficient and accurate grounding.
C.2 Why does CAPE enhance the performance?
The head-to-finger line provides a global directional prior, while the wrist-to-finger line captures local articulation. Together, they offer complementary spatial cues, allowing each network to capture different aspects of the pointing geometry and reducing the risk of misalignment across varied pointing poses. Therefore, we introduce CAPE module to ensemble the outputs of both models ( and ). We incorporate CLIP in addition to model confidence scores because CLIP learns a joint image–text embedding space where images and text with matching semantics are close. Formally, for an image region and a text description , CLIP provides a similarity score:
where and are image and text encoders, and is cosine similarity. This score captures semantic alignment.
Relying only on CLIP is insufficient because, while it provides semantic similarity, it does not encode geometric cues such as pointing direction. This can lead to imprecise localization in cluttered scenes or when multiple semantically plausible objects exist. Conversely, using only model confidence is also unreliable: pointing-based networks provide spatial priors but can be overconfident even when predictions are wrong due to occlusion, unusual poses, or network biases.
The two pointing models provide spatial priors, while CLIP provides semantic priors. Combining them reduces both geometric and linguistic ambiguity. Conceptually, the fusion functions like a Bayesian ensemble. The pointing models supply a geometry-driven likelihood over regions, while CLIP acts as a semantic prior. Integrating these signals produces more sharper and reliable posterior estimate.
As an alternative to CLIP model, we also experimented with SigLIP [80], which provides strong semantic grounding. However, replacing CLIP with SigLIP in the CAPE module yielded slightly lower performance (see Tab. C.8), so we continue using CLIP.
In summary, CAPE fuses pointing-based confidence with CLIP similarity, integrating complementary spatial and semantic signals to improve referent prediction in challenging poses.
| Method | IoU=0.25 | IoU=0.50 | IoU=0.75 |
| CAPE with SigLIP | 74.5 | 65.1 | 35.4 |
| CAPE with CLIP | 75.0 | 65.4 | 35.7 |
| Method | Heatmap Input | IoU=0.25 | IoU=0.5 | IoU=0.75 |
| Head-to-finger | 72.9 | 62.3 | 35.1 | |
| Wrist-to-finger | 73.0 | 62.0 | 35.6 | |
| Ground-Truth | 73.0 | 62.1 | 36.2 | |
| None | 71.6 | 60.7 | 32.9 | |
| Wrist-to-finger | 69.6 | 60.7 | 31.5 | |
| Head-to-finger | 68.1 | 59.2 | 28.4 | |
| Ground-Truth | 69.7 | 59.9 | 31.9 | |
| None | 68.8 | 59.7 | 31.2 |
C.3 Effect of Heatmap
We evaluate the robustness of our models ( and ) using different heatmap inputs (Table C.9). First, we test using a wrist-to-finger heatmap. While IoU 0.25 and 0.75 show slight gains, integrating this into CAPE reduces overall performance, indicating a loss of complementary effectiveness. In our main experiments, we generate both heatmaps using predicted keypoints from a pose estimation model. In contrast, the Ground-Truth setup uses the annotated pointing coordinates, including the center of the ground-truth bounding box, to generate the heatmap. As expected, this yields a small performance boost since the line is guaranteed to intersect the target object’s center. However, the improvement is marginal, indicating that our predicted heatmaps are already accurate and that the pointing line reliably guides the model. Finally, providing a zero heatmap (None) removes explicit pointing guidance. Performance drops but remains reasonable, demonstrating that the model can rely on text and visual context. Notably, since the model is trained with explicit pointing heatmaps, it has learned to interpret pointing cues; therefore, even without a heatmap at inference, it can still achieve meaningful performance. Repeating the same experiments for shows similar trends, reinforcing the complementary nature of the two heatmaps.
| Method | Text | IoU=0.25 | IoU=0.5 | IoU=0.75 |
| Original | 72.9 | 62.3 | 35.1 | |
| Dummy | 34.0 | 26.5 | 12.3 | |
| Random | 24.9 | 19.0 | 9.4 | |
| No Text | 39.2 | 27.9 | 10.2 | |
| Original | 69.6 | 60.7 | 31.5 | |
| Dummy | 34.6 | 25.5 | 11.0 | |
| Random | 25.4 | 19.3 | 9.24 | |
| No Text | 43.3 | 31.1 | 11.7 |
C.4 Effect of Text Input
In summary, for the embodied reference understanding task, text serves as the primary cue for object identification, while pointing information, derived from both the input image and the heatmap, plays a complementary, assistive role.
In Table C.10, we evaluate our models under varying text input conditions to assess the role of textual guidance. The Original setup uses the dataset annotations. In the Dummy scenario, the text is replaced with the generic word “object,” while in the Random scenario, text is randomly sampled from other test examples, sometimes contradicting the actual target. Both scenarios cause a notable performance drop, highlighting the importance of textual input. Without accurate text instructions, the pointing line often intersects multiple objects, making it insufficient to identify the correct referent in cluttered scenes.
In summary, text provides the primary cue for referent identification in ERU, while pointing information from the image and heatmap serves as a complementary signal.
C.5 Finetuning or Frozen Text Encoder?
We investigate the impact of freezing the text encoder during training. Based on multiple training experiments, we observed that finetuning the text encoder sometimes leads to instability, represented by the pink line in Fig. C.6. In contrast, freezing the text encoder results in consistently stable training, as shown by the green and orange lines in Fig. C.6. This indicates that finetuning introduces optimization sensitivity and may require multiple attempts to obtain a stable run. Despite this instability, finetuning ultimately yields better results. Table C.11 shows that, for both Setup A and Setup F, finetuning improves mAP at IoU 0.25 and 0.5, and also provides better scores. These gains suggest that adapting the text representations to the specific semantics of embodied reference understanding provides a meaningful advantage, even though the training process becomes less stable.
| Text Encoder | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Frozen, Setup A | 69.8 | 59.5 | 33.1 | 0.2454 | 0.2994 |
| Finetuned, Setup A | 71.2 | 60.1 | 32.8 | 0.2469 | 0.2662 |
| Frozen, Setup F | 71.3 | 62.1 | 33.7 | 0.2460 | 0.2634 |
| Finetuned, Setup F | 72.9 | 62.2 | 35.1 | 0.2457 | 0.2490 |
@@ C.5 Finetuning or Frozen Text Encoder?
Appendix D Hyperparameters and Other Training Parameters
In Tab. C.12, we present empirically determined hyperparameters and other key training details to support reproducibility. For most hyperparameters, we follow established values from the literature, as they have been thoroughly analyzed in prior work [38, 11, 37, 43, 81, 85, 39]. We also observe that training is not sensitive to the choice of loss coefficients, as the resulting accuracy differences are negligible.
Appendix E More Results
| IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | |||||||||
| Object Sizes | All | S | M | L | All | S | M | L | All | S | M | L |
| PaliGemma2 | 54.4 | 46.2 | 53.0 | 64.1 | 51.2 | 42.4 | 47.2 | 64.1 | 44.5 | 26.2 | 43.3 | 64.1 |
| Qwen2.5vl | 36.4 | 21.0 | 37.2 | 51.1 | 31.6 | 14.9 | 32.0 | 47.9 | 21.0 | 10.0 | 19.3 | 33.7 |
| Grounding DINO | 47.6 | 40.7 | 47.9 | 54.4 | 46.8 | 38.7 | 47.4 | 54.4 | 44.5 | 33.1 | 46.1 | 54.4 |
| Touch-Line-EWL | 50.7 | 49.8 | 53.2 | 45.6 | 46.5 | 45.5 | 49.3 | 41.3 | 24.6 | 20.3 | 30.9 | 29.3 |
| Touch-Line-VTL | 45.9 | 44.2 | 51.2 | 32.6 | 40.3 | 39.5 | 43.2 | 32.6 | 20.6 | 13.9 | 29.4 | 32.6 |
| 55.5 | 52.3 | 57.6 | 70.6 | 46.4 | 39.7 | 55.6 | 70.7 | 15.2 | 3.3 | 26.8 | 62.0 | |
| 48.6 | 46.2 | 49.4 | 64.1 | 42.5 | 40.1 | 43.1 | 59.8 | 17.7 | 7.8 | 28.8 | 43.5 | |
| CAPE w/ + | 55.5 | 52.9 | 55.8 | 75.0 | 50.5 | 47.3 | 50.9 | 75.0 | 18.6 | 9.1 | 26.6 | 57.6 |
In Tab. E.13, we report the results of our model on the CAESAR dataset, which is simulation-based. We compare our model with SOTA methods, including Touch-Line, GroundingDINO, PaliGemma2, and Qwen2.5vl. Despite the significant domain gap, our model achieves very accurate performance, attaining SOTA scores in most cases. Under the IoU threshold of 0.75, PaliGemma2 and GroundingDINO achieve higher scores.
In Tab. E.14, we report the mAP at three different thresholds for all ablation setups, categorized by object size. Similarly, in Tab. E.15, we present the CLIP scores and metrics for the same setups, also with respect to object size.
In Fig. E.7, we present example failure cases from the ISL Pointing and CAESAR datasets. In the first column, all models predict the same object, as it aligns better with the pointing line and the text provides limited semantic guidance. Localization-related text is particularly challenging for the models to interpret. In the second column, only Touch-Line-VTL predicts the correct object. In the third column (from ISL Pointing), our models detect the scissors correctly, whereas Touch-Line-VTL detects only a small part of the spoon. Due to perspective and the lack of depth information, our models align the scissors with the pointing line.
In Fig. E.8, we present additional example images from the YouRefIt dataset and qualitatively compare our model with SOTA models. CAPE achieves high accuracy in most cases, whereas Touch-Line often fails to detect the correct object, and the LMMs largely ignore the pointing cue, relying only on the text instruction.
In Fig. E.9, we present example images from the ISL Pointing and CAESAR datasets. The first two columns show ISL Pointing images, while the remaining two columns contain CAESAR samples, which are simulation-based. In the first ISL example, all models except predict the object correctly. makes wrong prediction because its prediction aligns better with the wrist-to-fingertip pointing line than the GT object, and the provided text does not conflict with this prediction. In the second example, fails, but CAPE correctly selects ’s prediction. Touch-Line-VTL, PaliGemma2, Qwen2.5vl, and GroundingDINO make mistakes as they overlook the pointing cue. On the CAESAR dataset, despite the severe domain gap, our model and Touch-Line-VTL achieve strong performance. Our models closely follow the pointing cue and detect the correct object more accurately than other models. Given the challenging text instructions in both datasets, accurately and efficiently leveraging the pointing direction is key, explaining why CAPE surpasses all other models.
@@ Appendix E More Results
| Touch-Line-EWL [38] | - | - | 0.2456 | 0.2312 | 0.2435 | 0.2615 | 0.3168 | 0.3006 | 0.2903 | 0.3564 | |
| Touch-Line-VTL [38] | - | - | 0.2456 | 0.2308 | 0.2440 | 0.2615 | 0.2843 | 0.2809 | 0.2276 | 0.3393 | |
| 1 | Baseline | - | - | 0.2470 | 0.2334 | 0.2447 | 0.2620 | 0.2662 | 0.2182 | 0.2506 | 0.3266 |
| 2 | Setup 1 + object center prediction | - | - | 0.2446 | 0.2316 | 0.2426 | 0.2596 | 0.2707 | 0.2201 | 0.2692 | 0.3222 |
| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 0.2451 | 0.2311 | 0.2432 | 0.2604 | 0.2984 | 0.2763 | 0.2616 | 0.3533 |
| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 0.2458 | 0.2312 | 0.2444 | 0.2617 | 0.2694 | 0.2649 | 0.2233 | 0.3172 |
| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 0.2448 | 0.2313 | 0.2435 | 0.2593 | 0.2770 | 0.2568 | 0.2319 | 0.3382 |
| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 0.2458 | 0.2318 | 0.2448 | 0.2606 | 0.249 | 0.222 | 0.2202 | 0.3024 |
| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 0.2448 | 0.2322 | 0.2426 | 0.2593 | 0.2744 | 0.2540 | 0.2314 | 0.3335 |
| 8 | Setup 6 | Channel-wise input | - | 0.2449 | 0.2311 | 0.2424 | 0.2609 | 0.2959 | 0.2437 | 0.2642 | 0.3767 |
| 9 | Setup 6 | Channel-wise feature | - | 0.2457 | 0.2318 | 0.2437 | 0.2613 | 0.2598 | 0.2289 | 0.2119 | 0.3339 |
| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 0.2578 | 0.2480 | 0.2614 | 0.2641 | 0.247 | 0.2176 | 0.2042 | 0.3141 |
| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 0.2644 | 0.2563 | 0.2670 | 0.2699 | 0.2391 | 0.2242 | 0.2011 | 0.2878 |
| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 0.2460 | 0.2656 | 0.2448 | 0.2606 | 0.2422 | 0.2158 | 0.2009 | 0.3060 |
| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 0.2581 | 0.2642 | 0.2485 | 0.2614 | 0.2434 | 0.2137 | 0.2028 | 0.3090 |
| 14 | Setup 5 + Setup 6 (Final Model) | Embedding feature | CAPE | 0.2661 | 0.2642 | 0.2672 | 0.2670 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 0.2461 | 0.2318 | 0.2441 | 0.2618 | 0.2634 | 0.2626 | 0.2233 | 0.3006 |
| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.2462 | 0.2321 | 0.2442 | 0.2620 | 0.2695 | 0.2685 | 0.2344 | 0.3028 |
| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.2464 | 0.2323 | 0.2444 | 0.2621 | 0.2765 | 0.2750 | 0.2499 | 0.3026 |
| Ours - Final | Embedding feature | CAPE | 0.2661 | 0.2642 | 0.2672 | 0.2670 | 0.2476 | 0.2137 | 0.2241 | 0.3023 | |